OpenStudy (anonymous):

A rectangular athletic field is twice as long as it is wide. Perimeter is 150 yards,what are the w and l?

OpenStudy (anonymous):



OpenStudy (anonymous):

im still lost

OpenStudy (anonymous):

Length = twice the width, right? so you could replace L by saying "2W" instead, right?

OpenStudy (anonymous):

Perimeter is the distance around the outside, so Length + Width + Length + Width or 2L + 2W if you simplify it. But you know that L is = 2W, because the problem says Length = twice the width. So P = 150 <---- given in problem 150 = 2L + 2W <<-- definition of perimeter 150 = 2(2W) + 2W <<--- replacing L with 2W 150 = 4W + 2W 150 = 6W W = 25 L = twice W = 2(25) = 50

OpenStudy (anonymous):

so l=65 ? right

OpenStudy (anonymous):

Um, no... Can you follow the steps I showed above? You start with knowing that perimeter is 150, so the sum of the 4 sides of the rectangle adds up to 150. The length sides are twice as long as the width sides, so you can add up perimeter by adding width + 2*width + width + 2*width... add that up to get perimeter is equal to 6*width. So 150 = 6*width, meaning width = 25. And Length is twice width, so twice 25 is 50. L is 50 Check work: Perimeter = 50 + 25 + 50 + 25 = 150 (right!) and length 50 is twice width 25 (right!!)